OpenStudy (anonymous):

Solve:ln2+lnx=5 a.50,00 b.74.2 c.10 d.3

OpenStudy (freckles):

ln(a)+ln(b)=ln(ab) ln(2)+ln(x)=?

OpenStudy (anonymous):

How do you solve..i need help

OpenStudy (freckles):

Ok do you see that one property of log I wrote: ln(a)+ln(b)=ln(ab) Try to use that here

OpenStudy (anonymous):

so plug in a and b

OpenStudy (freckles):

No I meant use this property that I have mentioned to rewrite ln(2)+ln(x)